OpenStudy (anonymous):

I have the answer can somone please check me... The equation of the line tangent to the curve y=x^3-6x^2 at x=2 is

OpenStudy (anonymous):

I think the answer is y=-12+8

OpenStudy (anonymous):

post answer choices plz

OpenStudy (anonymous):

choices a.y=-12+8 b. y=12x-12 c.y=-12x+40 d.y=12x-8 e.y=12x-40 f.y=-12+12

OpenStudy (anonymous):

it's a ur smart

OpenStudy (anonymous):

so choice A. is correct

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

-12x+8, but yeah :)

OpenStudy (anonymous):

wat